ENERGY TRANSFER COLLISIONAL PROCESS INVOLVING HETEROMOLECULAR COLLISIONS BE-TWEEN METHYL FLUORIDE AND N_{2}, Ar, He, CO_{2}, AND AIR

\underline{DANE J. PHILLIPS}, IERUS Technologies, 2904 Westcorp Blvd Ste 210, Huntsville, AL 35805; HENRY O. EVERITT, Army Aviation and Missile RD&E Center, Redstone Arsenal, AL 35898.

Time resolved IR/THz double resonance (DR) spectroscopy has been performed with a Q-switched CO_{2} laser and heterodyne THz detection. The rate constants associated with allowed rotational- and vibrational-state changing collisions of CH_{3}F with N_{2}, Ar, He, CO_{2}, and air are measured by monitoring the temporal evolution of the absorption strength for numerous rotational transitions as a function of pressure. Collision partner dependent energy transfer processes are studied and compared with homomolecular collisions. Energy transfer maps and associated collisional cross sections will be presented for each collision partner.